# Smoothed Graph Contrastive Learning via Seamless Proximity Integration

### v1

@@ A.2 Graph contrastive learning
To address this challenge, several Graph contrastive learning methods have emerged that eliminate the need for explicit negative pairs. For instance, BGRL applies the BYOL method (Grill et al., 2020) to graphs as a GCL approach that doesn’t depend on negative pairs (Thakoor et al., 2022). Additionally, Graph Barlow Twins (GBT) utilizes a cross-correlation-based loss function instead of negative samples (Bielak et al., 2022).
In our approach, we neither treat negative pairs the same way as in GRACE nor ignore them like in BGRL. Instead, we make use of negative pairs within the contrastive loss, but with a unique approach – we use the geometric structure of graphs to effectively consider proximity among negative pairs in contrastive learning, rather than treating them all the same.
Appendix B Mini-batch generating approaches
@@
Random edge sampler approach randomly selects edges from a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ based on a predefined probability distribution. For each edge $e$ in the set of edges $\mathcal{E}$, an independent decision is made to determine whether it should be included in the subgraph $\mathcal{G}_{s}$. This decision is guided by a probability value $P(e)$ assigned to each edge. The sampler incorporates a budget parameter $m$ to constrain the expected number of sampled edges, ensuring that $\sum P(e)=m$, as described in (Zeng et al., 2020).
Random walk sampler approach begins by randomly selecting $r$ root nodes as starting points on the entire graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. From each of these starting nodes, random walks of length $r$ are conducted to generate subgraphs (Zeng et al., 2020). To manage the potential issue of generating excessively large subgraphs, a batch size parameter $m$ is commonly employed, ensuring the approximate number of samples per batch.
Ego graph sampler approach generates subgraphs centered around a specific ”ego” node in a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. This mini-batch generation approach provides a localized perspective on the graph by constructing a $k$-hop ego-graph centered at node $v_{i}$, where ”$k$-hop” indicates that the subgraph includes nodes that can be reached within $k$ steps from $v_{i}$. Importantly, the sampler ensures that the maximum distance between $v_{i}$ and any other nodes within the ego-graph is limited to $k$, as expressed mathematically by $\forall v_{j}\in\mathcal{V},\mid d(v_{i},v_{j})\mid<k$ (Zhu et al., 2021a).
Appendix C Experimental setup
C.1 Properties and statistics of the benchmarks
For node classification, the benchmarks encompass a wide range of graph sizes, including smaller to medium-scaled ones such as Cora, Citeseer, Pubmed (Sen et al., 2008), CoauthorCs (Sinha et al., 2015), Computers, and Photos (McAuley et al., 2015), as well as larger datasets like ogbn-arxiv, ogbn-products, ogbn-proteins, and all of which are sourced from the Open Graph Benchmark (Hu et al., 2020). For graph classification, we employ MUTAG (Kriege & Mutzel, 2012), PTC (Kriege & Mutzel, 2012), IMDB-Binary (Yanardag & Vishwanathan, 2015), PROTEINS (Wale et al., 2008), and ENZYMES (Borgwardt et al., 2005) benchmarks.
@@ C.1 Properties and statistics of the benchmarks
Small Cora 2,485 5,069 1,433 7 78 80.4 4.08 19
Citeseer 2,120 3,679 3,703 6 438 73.5 3.47 28
[…]
@@
C.2 Hyperparameters
In all experiments, we follow the linear evaluation scheme outlined in (Veličković et al., 2019). Initially, we start by training the ’2-layer’ GCN encoders using the proposed SGCL framework in an unsupervised manner. The training process consists of 200 iterations, and we utilize the Adam optimizer with a learning rate of $31e-$ 3. Subsequently, the obtained embeddings are used to perform node or graph classification on a downstream task, employing a $l_{2}-$ regularized logistic regression classifier. The mean classification accuracy, along with the standard deviation, is then reported on the test nodes after conducting 5000 training runs.
In the mini-batch scenario of the node classification task, we employ a random-walk batch generation approach to create subgraphs from the input graph. We then conduct node classification for a downstream task, with a split ratio of 0.1/0.1/0.8 for train/validation/test. In both graph classification and the full-batch scenario of node classification, we follow the widely adopted data split used in self-supervised learning, as outlined in the Open Graph Benchmark.
To implement the proposed model, we leveraged the extensive capabilities offered by the PyGCL library, as introduced in (Zhu et al., 2021b). For the graph augmentation, we employ the augmentor base class provided by PyGCL, which includes Edge Removing (ER) and Node Feature Masking (FM), both with a drop probability of $0.5$.
Additionally, for a comprehensive comparison, we conducted experiments on all baselines using the PyGCL library since there was a lack of extensive baseline experimentation, especially for large-scale graphs.
In the mini-batch generation using the random-walk sampler, we set the batch size to 2000 with a random walk length of 4 and 3 starting root nodes for all benchmark datasets. However, for the ogbn-products benchmark, we use a batch size of 500 with a random walk length of 20.
In the smoothing techniques, we set the parameters as follows: For Taubin smoothing, we use $\mu=-0.4$, $\tau=0.3$, and $K=2$. In the case of Bilateral smoothing, we employ $\sigma_{spa}=0.1$ and $\sigma_{init}=2$. Finally, for Diffusion-based smoothing, we utilize $\eta=0.03$ and $K=2$.
@@ C.2 Hyperparameters
C.3 Baselines
In our empirical study, we incorporate a variety of models for comparison. For node classification, these models encompass representative node classification models, as well as recently-introduced graph contrastive learning models, such as DGI (Veličković et al., 2019), GRACE (Zhu et al., 2020), MVGRL (Hassani & Khasahmadi, 2020), GBT (Bielak et al., 2022), BGRL (Thakoor et al., 2022), CGRA (Duan et al., 2023), and GRLC (Peng et al., 2023) serving as our baseline models. For graph classification, we employ seven state-of-the-art methods for graph contrastive learning, including InfoGraph (Sun et al., 2019), GraphCL (You et al., 2020), MVGRL (Hassani & Khasahmadi, 2020), BGRL (Thakoor et al., 2022), AD-GCL (Suresh et al., 2021), LaGraph (Xie et al., 2022), and CGRA (Duan et al., 2023).
Appendix D Supplementary experiments
D.1 Node classification in full-batch scenario
Given the computational resource constraints, we are able to conduct experiments on small and medium-scale benchmarks using a full-batch scenario. As a result, to establish a fair and robust comparison with the existing state-of-the-art methods, we replicated the node classification experiment in a full-batch scenario, adhering to the commonly employed data split in self-supervised learning as provided in the Open Graph Benchmark. Table 6 demonstrates that, without any mini-batch generating step, SGCL outperforms state-of-the-art methods in four out of six benchmarks. However, it’s worth noting that the superiority of SGCL with the mini-batch generating step is more pronounced.
Table 6: Comparison of node classification accuracies of proposed models vs. baselines on small and medium-scaled graphs in full-batch scenario (mean ± std).
Model Cora Citeseer Pubmed CoauthorCS Computers Photo
DGI (Veličković et al., 2019) 76.28±0.04 69.33±0.14 83.79±0.08 91.63±0.08 71.96±0.06 75.27±0.02
GRACE (Zhu et al., 2020) 81.80±0.19 71.35±0.07 85.86±0.05 91.57±0.14 84.77±0.06 89.50±0.06
MVGRL (Hassani & Khasahmadi, 2020) 84.98±0.11 71.29±0.04 85.22±0.04 91.65±0.02 88.55±0.02 91.90±0.08
BGRL (Thakoor et al., 2022) 80.21±1.14 66.33±2.10 81.78±1.06 90.19±0.82 84.24±1.32 89.56±1.01
GBT (Bielak et al., 2022) 79.32±0.31 65.78±1.33 86.35±0.48 91.87±0.07 90.43±0.18 92.23±0.18
CGRA (Duan et al., 2023) 82.71±0.01 69.23±1.19 82.15±0.46 91.26±0.27 89.76±0.36 91.54±1.06
GRLC (Peng et al., 2023) 83.50±0.24 70.02±0.16 81.20±0.20 90.36±0.27 88.54±0.23 91.80±0.77

### v2

@@ A.2 Graph contrastive learning
Various models have explored different strategies to address the issue of negative pairs. For instance, BGRL applies the BYOL method [46] to graphs as a GCL approach that does not rely on negative pairs [21]. Similarly, Graph Barlow Twins (GBT) avoids the necessity for explicit negative pairs by utilizing a cross-correlation-based loss function [22].
Several studies focus on more informative negative samples, often referred to as hard negative samples. These samples closely resemble the anchor but have semantic differences. Intuitively, negative samples with different labels from the anchor, yet embedded nearby, are highly beneficial for providing significant gradient information during training. It’s preferable to choose negative pairs with very similar representations, as this makes it challenging for the current embedding to differentiate between them effectively. ProGCL [25] incorporates hard negative nodes into the contrastive loss to enhance performance. It applies a beta mixture model (BMM) to the pairwise similarities between the negatives and the anchor, estimating the probability of a negative being a true one. It subsequently integrates the estimated probability with the pairwise similarity to measure the hardness of the negative samples. AUGCL [27] is another hard negative mining GCL model that uses an affinity-based uncertainty estimator to evaluate the hardness of negative nodes relative to each anchor node. It constructs a discriminative model using pairwise affinities between negative nodes and the anchor, identifying nodes with higher uncertainty as hard negatives. Graph Soft-Contrastive Learning (GSCL) [47] is a novel approach aimed at overcoming the limitations of conventional graph contrastive learning by eliminating the need for graph augmentations and negative sampling. Instead, it leverages neighborhood ranking to ensure that closer nodes are more similar to a given anchor node than those farther away, in line with the inherent structure of the graph. The key limitation of GSCL is that it employs a specialized loss function for preserving the similarity ranking which requires computing a dense geodesic distance matrix for the entire graph. This process becomes increasingly challenging as the number of hops grows, leading to significantly higher computational costs, especially in large-scale graphs. Therefore, the relative similarity concept in GSCL is particularly well-suited to homophilic graphs, where label consistency decreases with distance.
In our approach, we neither treat negative pairs the same way as in GRACE nor ignore them like in BGRL or GSCL. Instead, we make use of negative pairs within the contrastive loss, but with a unique approach, we use the geometric structure of graphs to effectively consider proximity among negative pairs in contrastive learning, rather than treating them all the same. Integrating the proximity information to graph contrastive loss is still highly significant and to the best of our knowledge, our approach is the first work that addresses this limitation by promoting the geometric structure of data without encountering the limitations reported in hard negative mining methods. Specifically, our method overcomes the necessity of computing probability distributions and does not rely on prior assumptions, such as the bimodal similarity distribution of negatives with respect to positives as observed in ProGCL.
Appendix B Comprehensive overview of smoothing approaches
B.1 Taubin smoothing
Taubin smoothing is an iterative method that employs two distinct filters—positive and negative Laplacian filters—to enhance the smoothness of input data using the graph Laplacian matrix $\mathbf{L}$. Algorithm 1 outlines this process step-by-step.
The combination of positive and negative filters allows Taubin smoothing to balance effectively between smoothing and preserving features. It avoids excessive diffusion and oversmoothing by counteracting the smoothing effect with a corrective filter. Additionally, the parameters $\tau$ and $\mu$ enable fine control over the amount of smoothing and correction, making it adaptable to different graphs. However, it also has its limitations. Firstly, the method requires complex parameter tuning; balancing $\tau$ and $\mu$ effectively demands careful adjustment of hyperparameters, which can be challenging. Secondly, Taubin smoothing is sensitive to the underlying graph structure, heavily relying on the quality of the graph Laplacian. If the graph structure is irregular or contains noisy edges, the results may be adversely affected.
Algorithm 1 Taubin Smoothing $\mathcal{S}_{T}(\mathbf{V},\mathbf{L};K,\mu,\tau)$
1:  Input: Binary matrix $\mathbf{V}\in\{0,1\}^{N\times D}$, symmetric normalized graph Laplacian matrix $\mathbf{L}\in\mathbb{R}^{N\times N}$, number of iterations $K$, negative Laplacian filter constant $\mu(<0)$, positive Laplacian filter constant $\tau(>0)$ and $\mu<-\tau$
2:  Output: Smoothed matrix $\tilde{\mathbf{V}}$
3:  Initialize: Set $\mathbf{V}^{(0)}\leftarrow\mathbf{V}$
4:  Iterative Filtering:
5:  for $k\leftarrow 1$ to $K$ do
6:     Negative Laplacian Filter:
7:      Compute intermediate matrix $\mathbf{V}_{\text{temp}}^{(k)}$:
8:       $\mathbf{V}_{\text{temp}}^{(k)}\leftarrow(\mathbf{I}+\mu\mathbf{L})\mathbf{V}^{(k-1)}$
9:     Positive Laplacian Filter:
10:      Compute the updated matrix $\mathbf{V}^{(k)}$:
11:       $\mathbf{V}^{(k)}\leftarrow(\mathbf{I}+\tau\mathbf{L})\mathbf{V}_{\text{temp}}^{(k)}$
12:  end for
13:  for each node $i\textbf{ in }\{1,2,\ldots,N\}$
14:   If the original value $\mathbf{v}_{i}=1$, then
15:    Set $\tilde{\textbf{v}}_{i}\leftarrow 1$
16:  Return $\tilde{\mathbf{V}}$
B.2 Bilateral smoothing
This approach smooths input data by integrating spatial proximity and intensity similarity. This method considers both the distance between nodes in the graph and the similarity of their values to achieve effective smoothing. The intensity similarity $d_{\text{int}}(i,j)$ quantifies how similar two nodes are based on their intensity or binary values, with higher similarity indicating closer values. On the other hand, spatial proximity $d_{\text{spa}}(i,j)$ refers to the distance between two nodes $i$ and $j$ in the graph, measured by the shortest path. Algorithm 2 presents a step-by-step outline of the process.
Bilateral smoothing provides adaptive smoothing by responding to local differences in node values and spatial distances, enhancing its robustness in graphs with strong contrasts or noise. However, this approach can be computationally expensive on large graphs, as it requires calculating weights for each pair of nodes based on both spatial and intensity distances. Additionally, bilateral smoothing is sensitive to hyperparameters; the parameters $\sigma_{\text{spa}}$ and $\sigma_{\text{int}}$ must be carefully tuned, as improper values can lead to under- or over-smoothing.
Algorithm 2 Bilateral Smoothing $\mathcal{S}_{B}(\mathbf{V},\mathbf{A};\sigma_{\text{spa}},\sigma_{\text{int}})$
1:  Input: Binary matrix $\mathbf{V}\in\{0,1\}^{N\times D}$, adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, spatial smoothing parameter $\sigma_{\text{spa}}$, intensity smoothing parameter $\sigma_{\text{int}}$
2:  Output: Smoothed matrix $\tilde{\mathbf{V}}$
3:  for each pair of nodes $(i,j)$ in $\{1,2,\ldots,N\}$ where $j\in\mathcal{N}_{k}(i):$
4:   Compute weight $w(i,j)$:
5:    $w(i,j)\leftarrow\exp\left(-\frac{d_{\text{spa}}(i,j)}{2\sigma_{\text{spa}}^{2}}-\frac{d_{\text{int}}(i,j)}{2\sigma_{\text{int}}^{2}}\right)$
6:  for each node $i\textbf{ in }\{1,2,\ldots,N\}$
7:   Compute the smoothed value $\tilde{\textbf{v}}_{i}$:
8:    $\tilde{\textbf{v}}_{i}\leftarrow\frac{\sum_{j\in\mathcal{N}_{k}(i)}w(i,j)\textbf{v}_{j}}{\sum_{j\in\mathcal{N}_{k}(i)}w(i,j)}$
9:  for each node $i\textbf{ in }\{1,2,\ldots,N\}$
10:   If the original value $\mathbf{v}_{i}=1$, then
11:    Set $\tilde{\textbf{v}}_{i}\leftarrow 1$
12:  Return $\tilde{\mathbf{V}}$
B.3 Diffusion-based smoothing
Diffusion-based smoothing simulates the diffusion process (similar to heat diffusion) to propagate information across the graph. The value of each node diffuses into its neighbors, gradually smoothing the graph over time. Algorithm 3 presents a step-by-step outline of the process.
Diffusion-based smoothing is simple and efficient; the diffusion equation is relatively straightforward and computationally efficient to implement, making the method scalable for large graphs. Additionally, its iterative nature promotes smooth global effects, allowing values to propagate throughout the graph in a stable manner. However, there are notable disadvantages. One significant drawback is the potential loss of details; if not carefully controlled, the method can oversmooth the input, leading to the loss of sharp features or edges. Additionally, diffusion-based smoothing applies uniform smoothing; since it relies on averaging over neighboring nodes, it does not account for intensity similarity, which may result in the blurring of sharp changes in node values.
Algorithm 3 Diffusion-based Smoothing $\mathcal{S}_{D}(\mathbf{V},\mathbf{A};K,\eta)$
1:  Input: Binary matrix $\mathbf{V}\in\{0,1\}^{N\times D}$, adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, number of iterations $K$, diffusion rate $\eta$
2:  Output: Smoothed matrix $\tilde{\mathbf{V}}$
3:  Initialize $\mathbf{V}^{(0)}\leftarrow\mathbf{V}$
4:  for $k=0$ to $K-1$ do
5:     for each node $i\textbf{ in }\{1,2,\ldots,N\}$
6:      Compute the average value of neighboring nodes:
7:       $\bar{\textbf{v}}_{i}^{(k)}\leftarrow\sum_{j\in\mathcal{N}(v_{i})}\textbf{v}_{j}^{(k)}$
8:      Update the value of node $i$:
9:       $\textbf{v}_{i}^{(k+1)}\leftarrow\textbf{v}_{i}^{(k)}+\eta\bar{\textbf{v}}_{i}^{(k)}$
10:  end for
11:  for each node $i\textbf{ in }\{1,2,\ldots,N\}$
12:   If the original value $\mathbf{v}_{i}^{(0)}=1$, then
13:    Set $\textbf{v}_{i}^{(K)}\leftarrow 1$
14:  Return $\tilde{\mathbf{V}}\leftarrow\mathbf{V}^{(K)}$
In general, each smoothing approach has its unique strengths, making it appropriate for specific types of graph and smoothing requirements, allowing for tailored applications depending on the context. Taubin smoothing is ideal for graphs where connectivity is balanced and the local structure is not too irregular or noisy. Additionally, it is more efficient for large graphs as its computational complexity is lower compared to other methods in practice. Bilateral smoothing works well on graphs with heterogeneous or highly varying node values, such as those representing social networks, where sharp changes in node features are significant. Finally, Diffusion-based smoothing is best suited for large graphs with uniform or gradual changes, such as temperature distributions or geographical information, where computational efficiency is essential and the focus is on achieving smooth transitions without sharp features.
Appendix C Mini-batch generating approaches
@@
Random edge sampler approach randomly selects edges from a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ based on a predefined probability distribution. For each edge $e$ in the set of edges $\mathcal{E}$, an independent decision is made to determine whether it should be included in the subgraph $\mathcal{G}_{s}$. This decision is guided by a probability value $P(e)$ assigned to each edge. The sampler incorporates a budget parameter $m$ to constrain the expected number of sampled edges, ensuring that $\sum P(e)=m$, as described in [20].
Random walk sampler approach begins by randomly selecting $r$ root nodes as starting points on the entire graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. From each of these starting nodes, random walks of length $L$ are conducted to generate subgraphs [20]. To manage the potential issue of generating excessively large subgraphs, a batch size parameter $m$ is commonly employed, ensuring the approximate number of samples per batch.
Ego graph sampler approach generates subgraphs centered around a specific "ego" node in a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. This mini-batch generation approach provides a localized perspective on the graph by constructing a $k$-hop ego-graph centered at node $v_{i}$, where "$k$-hop" indicates that the subgraph includes nodes that can be reached within $k$ steps from $v_{i}$. Importantly, the sampler ensures that the maximum distance between $v_{i}$ and any other nodes within the ego-graph is limited to $k$, as expressed mathematically by $\forall v_{j}\in\mathcal{V},\mid d(v_{i},v_{j})\mid<k$ [48].
Appendix D Experimental setup
D.1 Properties and statistics of the benchmarks
For node classification, the benchmarks encompass a wide range of graph sizes, including smaller to medium-scaled ones such as Cora, Citeseer, Pubmed [49], CoauthorCs [50], Computers, and Photos [51], as well as larger datasets like ogbn-arxiv, ogbn-products, ogbn-proteins, and all of which are sourced from the Open Graph Benchmark [52]. For graph classification, we employ MUTAG [53], PTC [53], IMDB-Binary [54], PROTEINS [55], and ENZYMES [56] benchmarks.
@@ D.1 Properties and statistics of the benchmarks
Small Cora 2,708 5,429 1,433 7 78 80.4 4.08 19
Citeseer 3,327 4,732 3,703 6 438 73.5 3.47 28
Medium PubMed 19,717 44,324 500 3 1 80.2 4.5 18
@@
D.2 Hyperparameters
In all experiments, we follow the linear evaluation scheme outlined in [10]. Initially, we start by training the ’2-layer’ GCN encoders using the proposed SGCL framework in an unsupervised manner. The training process consists of 200 iterations, and we utilize the Adam optimizer with a learning rate of $31e-$ 3. Subsequently, the obtained embeddings are used to perform node or graph classification on a downstream task, employing a $l_{2}-$ regularized logistic regression classifier. The mean classification accuracy, along with the standard deviation, is then reported on the test nodes after conducting 5000 training runs.
In the mini-batch scenario of the node classification task, we employ a random-walk batch generation approach to create subgraphs from the input graph. We set the batch size to 2000 with a random walk length of 4 and 3 starting root nodes for all benchmark datasets. However, for the ogbn-products benchmark, we use a batch size of 500 with a random walk length of 20.
In the smoothing techniques, we set the parameters as follows: for Taubin smoothing, we set $\mu=-0.4$, $\tau=0.3$, and $K=2$; in the case of Bilateral smoothing, we employ $\sigma_{spa}=0.1$ and $\sigma_{init}=2$; and for Diffusion-based smoothing, we utilize $\eta=0.03$ and $K=2$.
To ensure a fair comparison with state-of-the-art models in both node and graph classification tasks, we adopt the widely used data split method from the Open Graph Benchmark, which is commonly employed in self-supervised learning. We also report values based on the respective papers. For benchmarks where experiments weren’t performed in the relevant papers, we accurately reproduced their values using available code resources. It’s worth mentioning that the results on the Cora in MVGRL [12] are reported across benchmarks with varying numbers of nodes and edges (refer to Table 1 in the respective paper). Therefore, the values are reproduced and reported using the standard Cora benchmark.
To implement the proposed model, we leveraged the extensive capabilities offered by the PyGCL library, as introduced in [57]. For the graph augmentation, we employ the augmentor base class provided by PyGCL, which includes Edge Removing (ER) and Node Feature Masking (FM), both with a drop probability of $0.5$. For a comprehensive comparison, we reported values based on the respective papers.
To ensure a fair comparison with state-of-the-art, we followed the publicly available data split of citation networks and replicated all experiments accordingly. For benchmarks where experiments weren’t performed in the relevant papers, we accurately reproduced their values using available code resources. Additionally, for large-scale graphs, we conducted experiments on most of the baselines using the PyGCL library since there was a lack of extensive baseline experimentation. The implementation is available at https://github.com/maysambehmanesh/SGCL.
@@ D.2 Hyperparameters
D.3 Baselines
In our empirical study, we incorporate a variety of models for comparison. For node classification, these models encompass representative node classification models, as well as recently-introduced graph contrastive learning models, such as DGI [10], GRACE [11], MVGRL [12], GBT [22], BGRL [21], CGRA [23], and GRLC [24] serving as our baseline models. For graph classification, we employ seven state-of-the-art methods for graph contrastive learning, including InfoGraph [29], GraphCL [30], MVGRL [12], BGRL [21], AD-GCL [31], LaGraph [32], and CGRA [23].
Appendix E Supplementary experiments
E.1 An empirical analysis of the feature space
In this section, we conduct the empirical analysis, introduced by Proposition 2.1, to validate the impact of the feature space on real-world graphs by calculating the disparity measure using two encoders within the proposed SGCL and conventional GCL (GRACE) frameworks. For a meaningful comparison of the disparity measures across all graphs, we normalized the values using Min-Max scaling, rescaling the measure to a range between 0 and 1. The results presented in Table 6 indicate that the mean disparity of the graph encoders across all variants of SGCL is consistently lower than the conventional GCL approach. This suggests that the SGCL encoder produces node representations with greater similarity within the same class and increased distinction between different classes, reflecting more effective self-supervised learning.
Table 6: Comparison of mean disparity measures for learned features of SGCL and GCL on real-world graph benchmarks.
Model Cora Citeseer Pubmed CoauthorCS Computers Photo ogbn-arxiv
GCL (GRACE) 0.66±0.05 0.63±0.02 0.51±0.03 0.58±0.02 0.64±0.03 0.65±0.03 0.56±0.03
SGCL-T 0.49±0.05 0.58±0.05 0.50±0.01 0.53±0.03 0.56±0.01 0.60±0.05 0.51±0.02
SGCL-B 0.63±0.03 0.51±0.05 0.49±0.02 0.52±0.01 0.52±0.05 0.59±0.02 0.48±0.05
SGCL-D 0.47±0.02 0.54±0.03 0.49±0.01 0.46±0.05 0.55±0.05 0.43±0.03 0.45±0.04
E.2 More evaluation on heterophilic graphs
To perform a more comprehensive analysis, we conduct experiments on graphs with varying homophily rates, utilizing different real-world graphs, as detailed in Table 7.